A cylindrical roller is of length 2m and diameter 84cm.find the number of revolutions it has to make to cover an area of 7920meter squared.
Answer by mananth(16949)

A cylindrical roller is of length 2m and diameter 84cm.find the number of revolutions it has to make to cover an area of 7920meter squared.
length of roller = 2m
radius of roller = 84/2 = 42 cm = 0.42 m
surface are of roller = 2pirl
=2*pi*0.42*2
=5.28 m^2
Number of revolutions = 7920/5.28
= 1500 revolutions